a block of mass 0.25Kg is placed on top of light vertical spring of force constant 500N/m pushed down so that the spring is compressed by 0.10m after the block is released from rest it travels upward and leaves the spring what is the maximum height


1
Expert's answer
2021-03-09T15:28:35-0500

Let's first find the velocity of the block after it released from the law of conservation of energy:


"PE=KE,""\dfrac{1}{2}kx^2=\dfrac{1}{2}mv^2,""v=\dfrac{kx^2}{m}=\dfrac{500\ \dfrac{N}{m}\cdot(0.1\ m)^2}{0.25\ kg}=20\ \dfrac{m}{s^2}."

Let's find the time that the block takes to reach the maximum height:


"v_y=v_{0}-gt,""0=v_{0}-gt""t=\dfrac{v_0}{g}=\dfrac{20\ \dfrac{m}{s}}{9.8\ \dfrac{m}{s^2}}=2.04\ s."

Finally, we can find the maximum height reached by the block:


"y_{max}=v_0t-\dfrac{1}{2}gt^2,""y_{max}=20\ \dfrac{m}{s}\cdot2.04\ s-\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(2.04\ s)^2=20.4\ m."